differentiate the following function with respect to x
x^(cos^-1 x)

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
take log both side gives
log y= log x(cos-1 x) = cos-1x logx
1/y * dy/dx = -1/sqrt(1-x^2) logx + cos-1 x/x
dy/dx= y(1/sqrt(1-x^2) logx + cos-1 x/x)